I've been trying to figure out the answer on the 3 variables equations but i can't find the solution. Supposibly the solution are whole numbers. Can you help me get the answer :
3a - 2b + 7c = 13
a + 8b - 6c = -47
7a - 9b -9c = -3

Answer by scott8148(6628)

multiplying 2nd eqn by -3 and adding to 1st eqn (eliminating "a") ___ -26b + 25c = 154 (A)

multiplying 2nd eqn by -7 and adding to 3rd eqn (eliminating "a") ___ -65b + 33c = 326 (B)

multiplying eqn (A) by -5 and adding to 2 times eqn (B) (eliminating "b") ___ -59c = -118

solve for c, and substitute back to find a and b